Daily report 8/30/02

Ted Berg tberg at snf.stanford.edu
Fri Aug 30 08:36:34 PDT 2002


Greetings all,
    The burn box was cleaned this morning, TCA saga continues, Leak rate
checked on BPSG and it is fine.


Date: 8/30/2002
Tube Name/Number Status Comments THICKNESS-Change at
Tylan 1 yellow Tube pulled and cleaned-Jim did steam ox investigating
TCA
Tylan 2 yellow TCA Investigation
Tylan 3 yellow TCA Investigation
Tylan 4 yellow TCA Investigation
Tylan 5-BBr3 Up
Tylan 6-POCl3 Up
Tylan 7-Dirty Ox Up
Tylan 8-SiGe up Waiting for feedback on Temp 5.2/???
Tylan 9-Poly Up  19.4/30
Tylan 10-Nitride Up  22.5/30
Tylan 11-BPSG-LTO Up  6.2/30
Tylan 12-FGA Up
Tystar Up  18.4/30???
Doped Poly Down
TEOS yellow TEOS Thickness OK -
Tylan 16 N/A
Burn Box Up Cleaned this morning

GAAS 22 UP
GAAS 23 UP

SiH4 for furnaces   Change at 100PSI Currently 500 PSI

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20020830/f94cac54/attachment.html>


More information about the diffusion mailing list